# Variability in the massive black hole binary candidate SDSS J2320+0024: no evidence for periodic modulation

### v1

@@
Variability in the massive black hole binary candidate SDSS J2320+0024: no evidence for periodic modulation
Fabio Rigamonti , fabio.rigamonti@inaf.it112244    Lorenzo Bertassi 33    Riccardo Buscicchio 3322    Fabiola Cocchiararo 3322    Stefano Covino 1144    Massimo Dotti 33 1 1 2 2    Alberto Sesana 33 1 1 2 2    and Paola Severgnini 11
@@
Massive black hole binaries (MBHBs) are a natural outcome of galaxy mergers, and they are expected to be among the loudest gravitational wave sources at low frequencies. SDSS J2320+0024 has been recently proposed as a promising MBHB candidate due to a possible periodicity in its light-curve and variability in the MgII emission line. In this work, we re-analyse the optical (g and r bands) light-curves of J2320+0024 within the framework of Bayesian model selection. When a periodicity is searched for together with red noise, the analysis of the g-band light-curve finds a peak in the posterior of the period at $\sim$290 days. The posterior profile is too broad to result in a preference for the periodic models with respect to models including only red-noise. Furthermore, the same peak is not present in the analysis of the r-band light-curve. A periodic model without red-noise identifies a different ($\sim$1100 days) periodicity, and it is significantly statistically disfavoured with respect to the other tested models. In summary, no significant evidence in favour of a true periodic signal over red noise variability is found. Our analysis questions the robustness of the previously claimed periodicity and emphasizes the importance of rigorous statistical treatment. While our findings challenge the binary interpretation for J2320+0024, they do not rule it out. A statistically robust joint analysis of the photometric light-curves and of the evolving broad line profiles can shed further light on the real nature of this object.
@@ 1 Introduction
In the current picture of hierarchical growth of galaxies, massive black hole (MBH) binaries (MBHBs) are expected to form and be fairly common in our Universe as a consequence of galaxy mergers (Begelman et al., 1980; Volonteri et al., 2016; Rosas-Guevara et al., 2019; De Rosa et al., 2019). The study, characterisation, and identification of these systems from an electromagnetic perspective are of paramount importance in light of current (e.g., pulsar timing array, PTA, Verbiest et al., 2016) and future (e.g., LISA Amaro-Seoane et al., 2017, 2022) gravitational waves (GWs) missions, as these systems are expected to be among the loudest sources of GWs. Still, MBHBs remain observationally elusive as several challenges complicate their unambiguous identification through electromagnetic signatures. In fact, to date, despite the large number of MBHB candidates being put forward, no definite observational confirmation of any of them has been provided, yet. The identification of MBHBs relies on indirect signatures either by searching for peculiar spectral features (Tsalmantza et al., 2011; Eracleous et al., 2012; Ju et al., 2013; Shen et al., 2013; Wang et al., 2017) or through photometric variability in their light-curve (Valtonen et al., 2008; Ackermann et al., 2015; Graham et al., 2015; Li et al., 2016; Charisi et al., 2016; Sandrinelli et al., 2016, 2018; Severgnini et al., 2018; Li et al., 2019; Liu et al., 2019; Chen et al., 2020; Serafinelli et al., 2020; Covino et al., 2019).
@@
The first approach is associated with MBHBs with separations larger than $0.01$ pc. At such distances, typically larger than the Roche lobe of the system, the individual MBHs still retain their own broad line region (BLR) (Montuori et al., 2011, 2012). In these cases, we expect broad emission lines (BELs) to be shifted in frequency with respect to their respective narrow emission lines (NELs) and to evolve in time over a binary orbital period. However, asymmetric emission line profiles can also be explained by complex dynamics and morphology of the BLR (Eracleous et al., 1997; Jovanović et al., 2010; Storchi-Bergmann et al., 2003; Rigamonti et al., 2025; Sottocorno et al., 2025). The presence of a SMBHB would be checked by measuring the expected drift in wavelengths over a binary orbital period, i.e. observing the candidate for $\approx 10{\rm yr}$ or longer  [^1] Notably, a faster test, requiring an observational campaign on timescales comparable to typical reverberation mapping observation, has been proposed by Gaskell (1988) and recently quantitatively detailed by Dotti et al. (2023)..
@@
The second approach is more suitable for binaries at closer separations, where the individual BLRs are either truncated or shared by both MBHBs. In such configurations, periodic variability over the orbital timescales is expected potentially due to periodic feeding from the circumbinary disc (Hayasaki et al., 2008; Tiede et al., 2024), Doppler boosted emission (D’Orazio et al., 2015), or periodic gravitational lensing (D’Orazio & Di Stefano, 2018; Davelaar & Haiman, 2022). However, even in that situations, convincing evidence that these sources are indeed MBHBs is still missing, both due the theoretical predictions for their emission and the plausibility of alternative interpretations such as the assumption of precession in jets and in the inner part of the accretion discs (Sandrinelli et al., 2016; Britzen et al., 2018). Solving this ambiguity requires alternative approaches, encompassing time domain and spectral analysis, (e.g., Bertassi et al., 2025) to robustly test the binary hypothesis. Finally, the most challenging complication in the identification of genuine MBHBs from light-curve analysis is given by the intrinsic quasar variability. Indeed the light-curves of accreting AGNs generally show correlated red-noise variability with power spectrum well described by a $\nu^{-2}$ power-law with a flattening at low frequency, i.e. a damped random walk (DRW, Kelly et al., 2009). Such variability can mimic quasi-periodicities, making actual periodicities difficult to unambiguously identify (Vaughan et al., 2016; Liu et al., 2018; Covino et al., 2019). Multiple objects initially showing an apparent periodicity were later revealed to be non-periodic (Graham et al., 2015; Liu et al., 2018; Jiang et al., 2022; Dotti et al., 2023). Established approaches to assess a source periodicity rely nowadays on Bayesian inference (Covino et al., 2020; Zhu & Thrane, 2020).
SDSS J2320+0024 (J2320+0024 hereinafter) has been proposed as a MBHB candidate due to a detected peak (T260similar-to-or-equalsabsent260$\simeq$ 260≃ 260 days) in its Lomb-Scargle periodogram (Fatović et al., 2023), possibly coupled with a variability in its Mg II broad emission line (Fatović et al., 2025). However, to date, no rigorous assessment of the significance of the observed periodicity has been provided. In particular, a comparison with quasar red noise and an evaluation within a robust Bayesian framework are still missing. In this study, we aim to test the binary nature of J2320+0024, focusing on the periodicity detected in its light-curve.
@@
In Sec. 2, we briefly present the archival datasets of the MBHB candidate analysed in this paper. We then describe our approach to the analysis of the light-curve of J2320+0024 in Sec. 3, ranking different noise and periodic models with each other. Finally, in Sec. 4 we discuss our results and present our conclusions.
@@ 2 Data
J2320+0024 is a system at redshift $z=1.05$, identified from a long-period variability search in the Sloan Digital Sky Survey Stripe 82 Standards Catalog (SDSS S82, York et al., 2000; Jiang et al., 2014; Ivezić et al., 2007). It was proposed as a MBHB candidate (Fatović et al., 2023) by Fatović et al. (2023), based on time-resolved analysis of its light-curve and, and was later followed up with spectral observation(Fatović et al., 2025). To assess the robustness of the periodicity detection in the light-curve of J2320+0024 we collected g- and r-band magnitudes from the SDSS, Pan-STARRS1 (PS1, Chambers et al., 2016), and Zwicky Transient Facility (ZTF, Graham et al., 2019; Bellm et al., 2019) surveys. Following Fatović et al. (2023), we calibrated the data from the different surveys to the median SDSS flux to empirically account for the instrumental offsets. In the following, we present the analysis on the g-band, while the corresponding results for the r-band are reported in Appendix A.
Table 1: Summary of the different model parameters for the fit to the g-band light-curve of SDSS J2320+0024.
@@
Scale length $\log_{10}{\tau}$ $[-3,1.7]$ $-1.3^{+0.3}_{-0.2}$ $-1.6^{+0.3}_{-0.2}$ $-1.3^{+0.4}_{-0.3}$
Dispersion $\log_{10}{\sigma}$ $[-3,1.7]$ $0.8^{+0.1}_{-0.1}$ $0.9^{+0.1}_{-0.1}$ $0.8^{+0.1}_{-0.1}$ $0.1^{+0.3}_{-0.2}$
Slope $\gamma$ $[0.5,5]$ $1.4^{+0.4}_{-0.4}$
Period $\log_{10}{T}$ $[-2,2]$ $0.2^{+1.3}_{-1.3}$ $-0.406^{+0.002}_{-0.002}$
@@
Scale length $\log_{10}{\tau}$ $[-3,1.7]$ $-1.35$ $-1.65$ $-0.74$
Dispersion $\log_{10}{\sigma}$ $[-3,1.7]$ $0.79$ $0.94$ $0.51$ $0.02$
Slope $\gamma$ $[0.5,5]$ $1.51$
Period $\log_{10}{T}$ $[-2,2]$ $-0.99$ $-0.41$
@@
 [^2] From left to right, the columns provide a brief description of each parameter, its reference name as used in this work, the assumed prior range, and the best-fit values with their credibility intervals for all the different models. The first table block reports the parameters estimated as the median, 16th, and 84th percentiles of the posterior distribution, while the second table block reports the set of parameters that maximise the likelihood. All the priors are log-uniform, and all the parameters are in dimensionless units since the data have been standardised before fitting. The last row reports the Bayesian log evidence of each model clearly disfavouring the Periodic model.
3 Modelling of the light-curve
@@
The same approach has already been proposed in the literature (Covino et al., 2020; Zhu & Thrane, 2020; Covino et al., 2022) and constitutes the most reliable solution to quasar light-curve characterization.
@@ 3.1 Gaussian Processes and Bayesian Inference
Bayesian analysis of light-curves typically makes use of the GP formalism (Rasmussen & Williams, 2006; Roberts et al., 2012; Angus et al., 2018). GPs describe distributions over functions and, in their discretized version, correspond to multivariate Gaussian distributions. GPs offer a flexible framework for modelling unknown time series by non-parametric models capable of capturing a large family of functions with two parameters: a kernel and a mean function. The kernel choice, which defines the GP covariance matrix, regulates the overall shapes and smoothness of samples, i.e. functions. Kernels can be of different functional form, usually depending on a few free (hyper)-parameters, and allowing for the modelling of either periodic or non-periodic light-curves. In the context of Bayesian inference, the identification of the kernel parameters that best describe the observed data passes through the evaluation of the GP log marginal-likelihood, which reads as follows Rasmussen & Williams (2006)
$\log{\mathcal{L}}(\mathbf{D}|\bm{\Omega})=-\frac{1}{2}(\mathbf{D}-\bm{\mu})^{T}\widetilde{{\bm{\Sigma}}}^{-1}(\mathbf{D}-\bm{\mu})-\frac{1}{2}\log{|\widetilde{\bm{\Sigma}}|}-\frac{N}{2}\log{2\pi},$ (1)
where, in our case, $\mathbf{D}$ is the observed magnitude, $\bm{\mu}$ is the mean function of the GP, $N$ is the number of observations, $\widetilde{\bm{\Sigma}}$ is the covariance matrix computed as the squared sum of the GP kernel $\bm{\Sigma}$ and the diagonal term given by the error measurements, while $|\widetilde{\bm{\Sigma}}|$ is the covariance matrix determinant. To reduce at minimum the number of free parameters in the model, we standardised (i.e., removed the mean and divided by the standard deviation [^3] The mean g-band magnitude is 21.616 while the standard deviation is 0.319.) the $\mathbf{D}$ so that we could always assume $\bm{\mu}=0$.
In this work, we compared four different kernels typically used to describe periodic and non-periodic light-curves: the exponential kernel DRW, the ”Modified DRW” kernel, the ”Quasi-periodic oscillation” (QPO) kernel, and the ”Periodic” kernel. We refer to Zhu & Thrane (2020) for a detailed discussion.
@@
where $\sigma$ is the intrinsic variance, $\tau$ is the damping timescale, and $t_{i}$ is the time at which the $i-$th observation has been taken. As done in the case of the magnitudes, also the observational times have been standardized [^4] The mean modified Julian date (MJD) is 56766.93, while the standard deviation is 2834.764.; this operation is done to reduce numerical issues when inverting the covariance matrix $\bm{\Sigma}$.
The DRW model can be generalized by including an additional free parameter $\gamma>0$ as an exponent to the $|t_{i}-t_{j}|/\tau$ term. This is done to provide an alternative that can account for quasar red noise deviating from the DRW model. This model, which we refer to as ”Modified DRW”, is characterized by the covariance matrix:
@@
Eqs. (2),(3) refer to pure noise models. When accounting for periodicity, kernels typically include terms proportional to an oscillating function. We define the quasi-periodic oscillation (QPO) kernel as
@@
containing a multiplicative periodic modulation with period $T$ to the DRW model. The QPO model accounts for a superposition of noise and a periodic signal and, in the limit of $T\to\infty$, reduces to the DRW model. In the absence of noise, the QPO kernel describes a purely oscillatory process
@@
which we refer to as ”Periodic” model. The final goal of this work is to measure the robustness of the presence of a periodic signal in the light-curve of J2320+0024. Bayesian statistics is perfectly suited for this task as it allows for precise model selection through the computation of the evidence of a model. The Bayesian evidence of a model ($Z$) is the normalization factor of the posterior distribution and, given the same set of data, allows for selecting the preferred among different models. Indeed, given two competing models $H_{i}$ and $H_{j}$, the Bayes ratio is:
$B_{ij}=\frac{p(D|H_{i})}{p(D|H_{j})}=\frac{Z_{i}}{Z_{j}}$ (6)
where we assume that there is no a priori preference for $H_{i}$ over $H_{j}$ and $Z_{i}/Z_{j}$. The ratio of the model evidences obtained by further marginalising over the entire parameter spaces for $H_{i}$ and $H_{j}$, is called the Bayes factor. If $B_{ij}$ is ¿ 1 (¡ 1) then model $H_{i}$ is favoured (disfavoured) compared to $H_{j}$. We refer to Jeffreys (1948) for a detailed classification and to Cocchiararo et al. (2025) for a more detailed discussion on GP inference in our astrophysical context.

### v2

# Variability in the supermassive black hole binary candidate SDSS J2320+0024: No evidence of periodic modulation

@@
Variability in the supermassive black hole binary candidate SDSS J2320+0024: No evidence of periodic modulation
Fabio Rigamonti , fabio.rigamonti@inaf.it    Lorenzo Bertassi    Riccardo Buscicchio    Fabiola Cocchiararo    Stefano Covino    Massimo Dotti    Alberto Sesana    and Paola Severgnini
@@
Supermassive black hole binaries ( SBHBs) are a natural outcome of galaxy mergers, and they are expected to be among the loudest gravitational-wave sources at low frequencies. The source SDSS J2320+0024 was recently proposed as a promising SBHB candidate due to a possible periodicity in its light curve and variability in the MgII emission line. In this work, we reanalysed the optical (g r, and i bands) light curves of J2320+0024 within the framework of Bayesian model selection. When periodicity was searched for together with red noise, analysis of the g-band light curve reveals a peak in the posterior of the period at $\sim$290 days. The posterior profile is too broad to yield a preference for periodic models over models that include only red noise. Furthermore, the same peak is not present in the analysis of the r-band and i-band light curve. A periodic model without red noise identified a different ($\sim$1100 days) periodicity, but this model is statistically significantly disfavoured relative to the other models tested. In summary, we find no significant evidence in favour of a true periodic signal over red-noise variability. Our analysis questions the robustness of the previously proposed periodicity and emphasises the importance of rigorous statistical treatment. While our findings challenge the binary interpretation for J2320+0024, they do not rule it out. A statistically robust joint analysis of the photometric light curves and evolving broad-line profiles would shed further light on the true nature of this object.
@@ 1 Introduction
In the current picture of hierarchical galaxy growth, supermassive black hole (SBH) binaries ( SBHBs) are expected to form and be fairly common in our Universe as a consequence of galaxy mergers (Begelman et al., 1980; Volonteri et al., 2016; Rosas-Guevara et al., 2019; De Rosa et al., 2019). The study, characterisation, and identification of these systems from an electromagnetic perspective is of paramount importance in light of current, e.g., pulsar timing array (PTA, Verbiest et al., 2016), and future, e.g., Laser Interferometer Space Antenna (LISA Amaro-Seoane et al., 2017, 2022) gravitational wave (GW) missions, as these systems are expected to be among the loudest sources of GWs. Nevertheless, SBHBs remain observationally elusive, as several challenges complicate their unambiguous identification through electromagnetic signatures. To date, despite the large number of SBHB candidates that have been put forward, no definite observational confirmation of any has yet been provided. The identification of SBHBs relies on indirect signatures, either through searches for distinctive spectral features (Tsalmantza et al., 2011; Eracleous et al., 2012; Ju et al., 2013; Shen et al., 2013; Wang et al., 2017) or through photometric variability in their light curve (Valtonen et al., 2008; Ackermann et al., 2015; Graham et al., 2015; Li et al., 2016; Charisi et al., 2016; Sandrinelli et al., 2016, 2018; Severgnini et al., 2018; Li et al., 2019; Liu et al., 2019; Chen et al., 2020; Serafinelli et al., 2020; Covino et al., 2019; Luo et al., 2025).
@@
The first approach is associated with SBHBs that have separations larger than $0.01$ pc. At such distances, typically larger than the Roche lobe of the system, individual SBHs still retain their own broad-line regions (BLRs) (Montuori et al., 2011, 2012). In these cases, we expect broad emission lines (BELs) to be shifted in frequency with respect to their narrow emission lines (NELs), evolving in time over a binary orbital period. However, asymmetric emission line profiles can also be explained by the complex dynamics and morphology of the BLR (Eracleous et al., 1997; Jovanović et al., 2010; Storchi-Bergmann et al., 2003; Rigamonti et al., 2025; Sottocorno et al., 2025). The presence of an SBHB can be tested by measuring the expected drift in wavelengths over a binary orbital period, i.e. by observing the candidate for $\approx 10{\rm yr}$ or longer  [^1] A faster test, requiring an observational campaign on timescales comparable to typical reverberation mapping observation, was proposed by Gaskell (1988) and more recently has been quantified by Dotti et al. (2023)..
@@
The second approach is more suitable for binaries at closer separations, where the individual BLRs are either truncated or shared by both black holes bound in the binary system. In such configurations, periodic variability on orbital timescales is expected, possibly due to periodic feeding from the circumbinary disc (Hayasaki et al., 2008; Tiede et al., 2024), Doppler-boosted emission (D’Orazio et al., 2015), or periodic gravitational lensing (D’Orazio & Di Stefano, 2018; Davelaar & Haiman, 2022). However, even in such situations, convincing evidence that these sources are indeed SBHBs is still missing, due both to theoretical predictions for their emission and to the plausibility of alternative interpretations, such as precession in jets and in the inner part of the accretion discs (Sandrinelli et al., 2016; Britzen et al., 2018). Solving this ambiguity requires alternative approaches, encompassing time-domain and spectral analyses, (e.g. Bertassi et al., 2025) to robustly test the binary hypothesis. Finally, the most challenging complication in the identification of genuine SBHBs from light-curve analysis arises from intrinsic quasar variability. Indeed, the light curves of accreting active galactic nuclei (AGNs) generally show correlated red-noise variability, with power spectra well described by a $\nu^{-2}$ power law that flattens at low frequency, i.e. a damped random walk (DRW; Kelly et al., 2009). Such variability can mimic quasi-periodicities, making genuine periodicities difficult to identify unambiguously(Vaughan et al., 2016; Liu et al., 2018; Covino et al., 2019). Multiple objects that initially appeared to show periodicity were later revealed to be non-periodic (Graham et al., 2015; Liu et al., 2018; Jiang et al., 2022; Dotti et al., 2023). Established approaches to assess source periodicity now rely on Bayesian inference (Covino et al., 2020; Zhu & Thrane, 2020; Hübner et al., 2022).
[…]
@@
In Sec. 2, we briefly present the archival datasets of the SBHB candidate analysed in this paper. We then describe our approach to analysing the light curve of J2320+0024 in Sec. 3, ranking different noise and periodic models against each other. Finally, in Sec. 4, we discuss our results and present our conclusions.
@@ 2 Data
J2320 + 0024 is a redshift system $z=1.05$, identified through a long-period variability search in the Sloan Digital Sky Survey Stripe 82 Standards Catalog (SDSS S82, York et al., 2000; Jiang et al., 2014; Ivezić et al., 2007; Thanjavur et al., 2021). It was proposed as an SBHB candidate by Fatović et al. (2023) based on time-resolved analysis of its light curve and, and was later followed up with spectral observations(Fatović et al., 2025). To assess the robustness of the periodicity detection in the light curve of J2320+0024, we collected g-, r-, and i-band magnitudes from the SDSS, Pan-STARRS1 (PS1, Chambers et al., 2016), and Zwicky Transient Facility (ZTF, Graham et al., 2019; Bellm et al., 2019) surveys. Following Fatović et al. (2023), we calibrated the data from the different surveys with the median SDSS flux to empirically account for instrumental offsets. In the following, we present the analysis on the g-band, while the corresponding results for the r- and i-band are reported in Appendix A and Appendix B respectively.
Table 1: Summary of model parameters for the g-band light-curve fit of SDSS J2320+0024.  [^2] From left to right, the columns provide a brief description of each parameter, its reference name as used in this work, the assumed prior range, and the best-fit values with their credibility intervals for all the models. The first table block reports parameters estimated as the median, 16th, and 84th percentiles of the posterior distribution, while the second table block reports the parameter set that maximises the likelihood. All priors are log-uniform and all the parameters are in dimensionless units since the data were standardised before fitting. The last row reports the Bayesian log evidence of each model, clearly disfavouring the Periodic model.
@@
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] $1.3^{+0.3}_{-0.2}$ $1.6^{+0.3}_{-0.2}$ $1.3^{+0.4}_{-0.3} -$
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.8^{+0.1}_{-0.1}$ $0.9^{+0.1}_{-0.1}$ $0.8^{+0.1}_{-0.1}$ $0.1^{+0.3}_{-0.2}$
Slope $\gamma$ $[0.5,5]$ - $1.4^{+0.4}_{-0.4}$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - $0.2^{+1.3}_{-1.3}$ $0.406^{+0.002}_{-0.002}$
@@
Scale length $\log_{10}{\tau}$ [3,1.7][-3,1.7] 1.35-1.35 1.65-1.65 0.74-0.74
Dispersion $\log_{10}{\sigma}$ [3,1.7][-3,1.7] $0.79$ $0.94$ $0.51$ $0.02$
Slope $\gamma$ $[0.5,5]$ - $1.51$ - -
Period $\log_{10}{T}$ [2,2][-2,2] - - 0.99-0.99 0.41-0.41
@@
3 Modelling of the light curve
@@
This approach has been proposed previously in the literature (Covino et al., 2020; Zhu & Thrane, 2020; Covino et al., 2022; Hübner et al., 2022) and constitutes the most reliable method for quasar light-curve characterisation.
@@ 3.1 Gaussian processes and Bayesian inference
Bayesian analysis of light curves typically makes use of the GP formalism (Rasmussen & Williams, 2006; Roberts et al., 2012; Angus et al., 2018). Gaussian processes describe distributions over functions and, in their discretised version, correspond to multivariate Gaussian distributions. This framework offers a flexible approach for modelling unknown time series through non-parametric models capable of capturing a large family of functions with two parameters: a kernel and a mean function. The kernel choice, which defines the GP covariance matrix, regulates the overall shapes and smoothness of samples, i.e. functions. Kernels can have different functional forms, usually depending on a few free (hyper)-parameters, allowing the modelling of either periodic or non-periodic light curves. In Bayesian inference, identifying the kernel parameters that best describe the observed data requires evaluating the GP log marginal likelihood, which is given by Rasmussen & Williams (2006)
$\log{\mathcal{L}}(\mathbf{D}|\boldsymbol{\Omega})=-\frac{1}{2}(\mathbf{D}-\boldsymbol{\mu})^{T}\widetilde{{\boldsymbol{\Sigma}}}^{-1}(\mathbf{D}-\boldsymbol{\mu})-\frac{1}{2}\log{|\widetilde{\boldsymbol{\Sigma}}|}-\frac{N}{2}\log{2\pi}.$ (1)
[…]
We compared four different kernels typically used to describe periodic and non-periodic light curves: the exponential kernel DRW, the modified DRW kernel, the QPO kernel , i.e. a superposition of noise and a periodic signal, and the periodic kernel. We refer the reader to Zhu & Thrane (2020) for a detailed discussion.
@@
where $\sigma$ is the intrinsic variance, $\tau$ is the damping timescale, and $t_{i}$ is the time at which the $i$-th observation was taken. As with the magnitudes, the observational times were also standardised [^4] The mean modified Julian date (MJD) is 56766.93, while the standard deviation is 2834.764.. This reduced numerical issues when inverting the covariance matrix $\boldsymbol{\Sigma}$.
The DRW model can be generalised by including an additional free parameter $\gamma>0$ as an exponent to the $|t_{i}-t_{j}|/\tau$ term. This modification accounts for quasar red noise that deviates from the DRW model. This model, referred to as the modified DRW, is characterised by the covariance matrix
@@
Eqs. (2),(3) represent pure noise models. When accounting for periodicity, kernels typically include terms proportional to an oscillating function. We define the quasi-periodic oscillation (QPO) kernel as
@@
which contains a multiplicative periodic modulation with period $T$ to the DRW model. The QPO model accounts for a superposition of noise and a periodic signal and reduces to the DRW model in the limit $T\to\infty$. Gaussian processes are a flexible, intrinsically stochastic approach to model functions. The shape of the covariance matrix regulates the smoothness of the GP realisations but is not reflected in a unique shape of the output function (i.e. it is not deterministic). For example, the covariance in Eq. 4 does not restrict the model to purely cosine-like periodicities, but allows for more general signals with probabilistic periodic correlations. In the absence of noise, the QPO kernel describes a purely oscillatory process:
@@
which we refer to as the periodic model. The final goal of this work was to measure the robustness of the presence of a periodic signal in the light curve of J2320+0024. Bayesian statistics is well suited for this task, as it allows precise model selection through the computation of the model evidence. The Bayesian evidence of a model ($Z$) is the normalisation factor of the posterior distribution and enables the selection of the preferred model given the same dataset. Given two competing models, $H_{i}$ and $H_{j}$, the Bayes ratio is
$B_{ij}=\frac{p(D|H_{i})}{p(D|H_{j})}=\frac{Z_{i}}{Z_{j}},$ (6)
where no a priori preference for $H_{i}$ over $H_{j}$ and $Z_{i}/Z_{j}$ is assumed. The ratio of the model evidences obtained by further marginalising over the entire parameter spaces for $H_{i}$ and $H_{j}$ is called the Bayes factor. If $B_{ij}$ is ¿ 1 (¡ 1), then model $H_{i}$ is favoured (disfavoured) compared to $H_{j}$. We refer to Jeffreys (1948) for a detailed classification and to Cocchiararo et al. (in prep.) for a more detailed discussion of GP inference in our astrophysical context.